# Unbiased Primordial Gravitational Wave Inference from the CMB with SMICA

### v2

@@ 2 Mock Observations
In this work we consider instrument parameters and science targets in line with those of CMB-S4. We define a possible sky patch that would be similar to the CMB-S4 ‘ultra-deep’ sky patch [10], with a noise model adapted from the most recently published CMB-S4 forecasting paper [21], as detailed in Table 1. The various frequency channels are selected for discerning between CMB signals, the frequency dependence of which are those of the temperature derivative of a 2.726 K blackbody, and those of Galactic foreground emissions. We generate 96 different noise realizations, and an equal number of CMB realizations, for a statistically significant sample of mock observations to test the pipeline.
@@ 2.1 Sky Patch
where $w_{p}$ is the window function that describes an apodized mask shown in Figure 1.
@@ 2.2 Noise Model
where $\Delta_{P}$ is the white noise level, and $\ell_{\rm knee}$ and $\alpha_{\rm knee}$ specify the low-$\ell$ contribution from the projected $1/f$ noise. The values of these three noise parameters are frequency dependent. The parameter values used in our simulation are listed in Table 1, and are based on the the most recently published CMB-S4 forecasting paper [21].
@@
Figure 1: Sky patch (left) with binary outline traced in black, centered at (RA = $10^{\circ}$, dec = $45^{\circ}$). The apodization yields $f_{\text{sky}}=2.5\%$ as calculated with Eq. (2.1). The sky map is polarized intensity of the PySM3 high-complexity foregrounds at 270 GHz. Plot of beam deconvolved noise curves (right) from noise levels in Table 1 for both experimental configurations. Overlaid are the theoretical CMB signals – lensing signal (black) and primordial $B$-mode signal (red). The $\ell$-range used in the analysis is bounded by the dotted vertical black lines.
@@ 2.3 Sky Models
In our simulation pipeline we use the recommended foreground model suites from the Python Sky Model 3 (PySM3) [22, 23, 24]. In this work we only include the polarized foreground components, ignore all extragalactic foregrounds, and use mono-frequency bandpasses, so we also ignore CO line-emissions. The large scale modes of the dust models from PySM3 are based on the Planck GNILC dust map, while the polarized synchrotron are based on the WMAP K-band observations. The small scales are generated from best-fit power law model of the the dust/synchrotron power spectra. The low-complexity foregrounds (synchrotron s4, dust d9) uses templates produced as described above, and their frequency scaling is constant over the entire sky. The medium-complexity model (synchrotron s5, dust d10) uses the exact same spatial template as the low-complexity, with spatial variations in the frequency scaling. The high-complexity foregrounds (synchrotron s7, dust d12, anomalous microwave emission a2) include an additional polarized foreground in form of the anomalous microwave emission (AME) based on the Planck Commander AME data product. The synchrotron model introduces spectral running in the power law frequency scaling, while the dust is a multi-layer model with line-of-sight decorrelation generated following the prescription of [25].
@@ 3.2 Implementation
The SMICA data covariance model of Eq. (3.1), (negative) log-likelihood of Eq. (3.5), and the sampling setup form the core of the SMICA pipeline. A flowchart for this pipeline is shown in Figure 3. The covariance model and likelihood are implemented in JAX [31], making use of JAX’s just-in-time compilation and vectorized-mapping features, allowing for fast computations. The log-likelihood is sampled with a No-U-Turn Sampler (NUTS) which pairs gradient descent with Monte-Carlo sampling. The NUTS sampling is implemented with BlackJax [32], and makes use of JAX’s automatic differentiation features for computing the gradients of the log-likelihood. The automatic differentiation also provides fast Fisher forecasts, which we use to inform our initial conditions. The choice of $n_{\rm FG}$ for the model is informed from the noise-whitened observation covariance matrix similar to [33], and is described in Sec A.
@@ 4 Results
We validate our pipeline with the low-complexity foregrounds which have only two rigid-scaling emission components; therefore we only need two foreground components to obtain an unbiased fit for $r$. Figure 4 shows the posterior probability distribution (normalized at its peak). Tables 3 and 3 show the MAP estimate of $r$, the uncertainty, $\sigma_{r}$, (computed as the standard deviation of the MC chains) and the $\chi^{2}/n_{\text{dof}}$ goodness-of-fit metric for $r=0$ and $r=3\times 10^{-3}$. Note that the expectation value for $r$ (minimum mean squared error estimate) is nearly identical to the MAP estimate. We obtain an unbiased fit for $r$ for both instrument configurations and the standard deviation $\sigma_{r}$ is small, exceeding the CMB-S4 $r$-sensitivity target. Additionally, the standard deviation obtained matches the Fisher forecast for $\sigma_{r}$ which validates our sampling method. The $\chi^{2}/n_{\text{dof}}$ metric is close to one, indicating a reasonably good fit. When we attempt this fit to low-complexity foregrounds with $n_{\text{FG}}>2$ the simulation fails to converge; BlackJAX appears to fail when overfitting data.
@@
While $\sigma_{r}$ increases with the number of parameters, the $\chi^{2}/n_{\text{dof}}$ generally improves with additional components in our model as is indicated by Tables 3 and 3. It is important to note that $\chi^{2}/n_{\text{dof}}$ is not far from one anywhere, even in the cases with significant $r$-bias. This is most likely because the relative contribution of the primordial $B$-mode to the data covariance model is vanishingly small. The foreground residuals in the CMB frequencies are consistent with zero within $1\sigma$ in all models and configurations, as shown in appendix B. This implies that the $\chi^{2}/n_{\text{dof}}$ is not very useful when it comes to informing us about the $r$-bias. We do not find any $\chi^{2}/n_{\text{dof}}$ metric smaller than 1, so even with the additional parameters there is no indication that we are over-fitting the system.

### v3

@@ 2 Mock Observations
In this work we consider instrument parameters and science targets in line with those of CMB-S4. We define a possible sky patch that would be similar to the CMB-S4 ‘ultra-deep’ sky patch [1], with a noise model adapted from the most recently published CMB-S4 forecasting paper [5], as detailed in Table 1. Two options are considered: One of them, called “non-split”, has a total of 7 frequency bands, with two main CMB channels at 90 and 150 GHz, three low frequency channels, and two high frequency channels. The other, labelled as “split”, replaces the two CMB channels by four channels centered at 85, 95, 145, and 155 GHz, for a total of 9 bands. The various frequency channels are selected for discerning between CMB signals, the frequency dependence of which are those of the temperature derivative of a 2.726 K blackbody, and those of Galactic foreground emissions. We generate 96 different noise realizations, and an equal number of CMB realizations, for a statistically significant sample of mock observations to test the pipeline.
We note that the map-based simulations used in this work do not capture all possible complexities of the data properties of ground-based CMB measurements. Processing of the data timestreams with various filters and template subtractions would produce effects and residual systematics that are not captured by our simulation pipeline. We postpone the inclusion of such effects and systematics to a future project.
@@ 2.1 Sky Patch
where $w_{p}$ is the window function that describes an apodized mask shown in Figure 1. For the purpose of this work we assume that all instruments are observing the same patch from the same site at the South Pole.
@@ 2.2 Noise Model
where $\Delta_{P}$ is the white noise level, and $\ell_{\rm knee}$ and $\alpha_{\rm knee}$ specify the low-$\ell$ contribution from the projected $1/f$ noise. The values of these three noise parameters are frequency dependent. The parameter values used in our simulation are listed in Table 1, and are based on the most recently published CMB-S4 forecasting paper [5].
@@
Figure 1: Sky patch (left) with binary outline traced in black, centered at (RA = $10^{\circ}$, dec = $45^{\circ}$). The apodization yields $f_{\text{sky}}=2.5\%$ as calculated with Eq. (2.1). The sky map is polarized intensity of the PySM3 high-complexity foregrounds at 270 GHz. Plot of beam deconvolved noise curves (right) from noise levels in Table 1 for both experimental configurations. We also plot the expected foregrounds from PySM (shaded areas) at select frequencies (40, 90, 150, and 220 GHz) where the width is given by the discrepancy between the low, medium, and high complexity models. Overlaid are the theoretical CMB signals – lensing signal (black) and primordial $B$-mode signal (red). The $\ell$-range used in the analysis is bounded by the dotted vertical black lines.
@@ 2.3 Sky Models
In our simulation pipeline we use the recommended foreground model suites from the Python Sky Model 3 (PySM3) [32, 34, 31]. In this work we only include the polarized foreground components, ignore all extragalactic foregrounds, and use monochromatic bandpasses, so we also ignore CO line-emissions. The large scale modes of the dust models from PySM3 are based on the Planck GNILC dust map, while the polarized synchrotron are based on the WMAP K-band observations. The small scales are generated from best-fit power law model of the dust/synchrotron power spectra. The templates are defined at 353 GHz for the dust and 23 GHz for the synchrotron. The foregrounds at other frequencies are produced by rescaling these templates to those frequencies assuming a modified blackbody emission law for the dust or a power law (with or without spectral running) emission for the synchrotron.
The low-complexity foregrounds (PySM3 synchrotron model s4, PySM3 dust model d9) uses templates produced as described above, and their frequency scaling is constant over the entire sky. The medium-complexity model (PySM3 synchrotron model s5, PySM3 dust model d10) uses the exact same spatial template as the low-complexity, with spatial variations in the frequency scaling. The high-complexity foregrounds (PySM3 synchrotron model s7, PySM3 dust model d12, PySM3 anomalous microwave emission model a2) include an additional polarized foreground in form of the anomalous microwave emission (AME) based on the Planck Commander AME data product. The synchrotron model introduces a spectral running in the power law frequency scaling, while the dust is a multi-layer model with line-of-sight decorrelation generated following the prescription of [24]. We refer the reader to tables 3, 4 and 7 of [31] for a tabulated summary of the PySM3 foreground models. These recommended PySM3 foreground suites span the current range of foreground modeling uncertainty, and the power spectra of select frequencies are plotted in Figure 1. Going through the model suites from low to high, increases the number polarized foreground components and frequency decorrelation for the dust and synchrotron. This makes the component separation problem more difficult as we go from low to medium to high complexity foregrounds.
@@ 3.2 Implementation
The SMICA data covariance model of Eq. (3.1), (negative) log-likelihood of Eq. (3.5), and the sampling setup form the core of the SMICA pipeline. A flowchart for this pipeline is shown in Figure 3. The covariance model and likelihood are implemented in JAX [13], making use of JAX’s just-in-time compilation and vectorized-mapping features, allowing for fast computations. The log-likelihood is sampled with a No-U-Turn Sampler (NUTS) which pairs gradient descent with Monte-Carlo (MC) sampling. The NUTS sampling is implemented with BlackJax [6], and makes use of JAX’s automatic differentiation features for computing the gradients of the log-likelihood. The automatic differentiation also provides fast Fisher forecasts, which we use to inform our initial conditions. The choice of $n_{\rm FG}$ for the model is informed from the noise-whitened observation covariance matrix similar to [29], and is described in Sec A.
@@ 4 Results
We validate our pipeline with the low-complexity foregrounds which have only two rigid-scaling emission components; therefore we only need two foreground components to obtain an unbiased fit for $r$. Figure 4 shows the posterior probability distribution (normalized at its peak). Tables 3 and 3 show the MAP estimate of $r$, the uncertainty, $\sigma_{r}$, (computed as the standard deviation of the MC chains) and the $\chi^{2}/n_{\text{dof}}$ goodness-of-fit metric for $r=0$ and $r=3\times 10^{-3}$. Note that the expectation value for $r$ (minimum mean squared error estimate) is nearly identical to the MAP estimate. We obtain an unbiased fit for $r$ for both instrument configurations and the standard deviation $\sigma_{r}$ is small, exceeding the CMB-S4 $r$-sensitivity target. Additionally, the standard deviation obtained matches the Fisher forecast for $\sigma_{r}$ which validates our sampling method. The $\chi^{2}/n_{\text{dof}}$ metric is close to one, indicating a reasonably good fit. When we attempt this fit to low-complexity foregrounds with $n_{\text{FG}}>2$, we cause a degeneracy in the foreground parameters. When we fit for 3 or 4 foreground components instead of 2, instabilities are observed in the MC chains for the foreground parameters specifically. We also find the $\chi^{2}/n_{\text{dof}}$ increases because we are adding parameters to the model without obtaining a better fit to the data. Though we have degeneracies and lack of convergence in the foregrounds, the posterior of the $r$ parameter specifically remains well constrained and remains largely unchanged, showing that the method is robust.
@@
While $\sigma_{r}$ increases with the number of parameters, the $\chi^{2}/n_{\text{dof}}$ generally improves with additional components in our model as is indicated by Tables 3 and 3 up to $n_{\text{FG}}=4$ for both medium and high complexity foregrounds. It is important to note that $\chi^{2}/n_{\text{dof}}$ is not far from one anywhere, even in the cases with significant $r$-bias. This is most likely because the relative contribution of the primordial $B$-mode to the data covariance model is vanishingly small. The foreground residuals in the CMB frequencies are consistent with zero within $1\sigma$ in all models and configurations, as shown in appendix B. This implies that the $\chi^{2}/n_{\text{dof}}$ is not very useful when it comes to informing us about the $r$-bias. We do not find any $\chi^{2}/n_{\text{dof}}$ metric smaller than 1, so even with the additional parameters there is no indication that we are over-fitting the system.
@@ 5 Discussion
We note that as is the case for any fitting procedure, SMICA can only work if the model that is fit is not wildly underconstrained, leading degeneracies in the parameters of interest. Still, we find convergence to the correct value of $r$ even in cases for which only the foreground parameters suffer from degeneracies, as is the case when we try to fit for 3 or 4 foreground components in the low-complexity case.